# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

### v3

@@ D.1 Simulation design
We generate data from a discrete-state Markov process, where each individual is characterized by a state $S=(\text{engagement},\text{churn risk},\text{tenure},\text{overlap})$, with each variable taking values in $\{0,1,2\}$. The initial state follows $\text{engagement}\sim\text{Multinom}(0.5,0.3,0.2)$, $\text{churn risk}\sim\text{Multinom}(0.25,0.25,0.25)$, $\text{tenure}\sim\text{Multinom}(0.25,0.25,0.25)$, and $\text{overlap}\sim\text{Multinom}(0.7,0.3,0.2)$. Treatment is assigned as $Z\sim\text{Bernoulli}(\pi)$ with $\pi=0.25$. State transitions evolve as follows: tenure increments deterministically as $T_{t+1}=\min(T_{t}+1,2)$. Engagement follows a random walk, where the probability of decrementing is $p_{0}(s_{t})=0.8-C_{t}/5$ for $Z=0$ and $\min(0.1+(0.8-C_{t}/5),1)$ for $Z=1$, with $E_{t+1}=\min(\max(E_{t}+2B_{0}-1,0),2)$, where $B_{0}\sim\text{Bernoulli}(p_{0}(Z))$. Churn risk evolves similarly, with $p_{1}(Z)=0.6$ for $Z=0$ and $0.4$ for $Z=1$, and updates as $C_{t+1}=\min(\max(C_{t}+2B_{1}-1,0),2)$, where $B_{1}\sim\text{Bernoulli}(p_{1}(Z))$. Overlap updates as $O_{t+1}=\min(O_{t}+1,2)$ if $Z=1$ and $B_{2}=1$, otherwise $O_{t+1}=0$, where $B_{2}\sim\text{Bernoulli}(\beta)$. The overlap parameter $\beta$ controls the degree of overlap between states over time, with larger values indicating less overlap. The reward is generated as $Y_{t}\mid(S_{t},Z)\sim\text{Bernoulli}(\sigma(-0.5+1\{O_{t}>0\}+T_{t}/2+0.3Z+1\{E_{t}>0\}/2-C_{t}/2))$, where $\sigma(x)=1/(1+e^{-x})$.
@@ D.2 Additional experimental results
Appendix E Details on data-fusion application
E.1 Background on confounding-robust data-fusion
We consider a data fusion setting in which experimental data ($Z=1$) is augmented with historical control data ($Z=0$), adapting the frameworks of Kallus et al. (2018) and van der Laan et al. (2024d) to Markov decision processes. Define the state as $S_{t}=(Z,\widetilde{S}_{t})$, as in Example 2. Suppose the study indicator $Z$ denotes whether a study unit belongs to an observational study $(Z=0)$, such as historical data, or a randomized experiment $(Z=1)$, such as a randomized control trial. We address the data-fusion problem of augmenting experimental data with potentially biased observational data to increase power by effectively enlarging the sample size. While randomization ensures unbiasedness in the experimental study, incorporating observational data can introduce bias from unmeasured confounding unless strong, untestable assumptions are made. Our goal is to combine these data sources in a confounding-robust manner that retains the unbiasedness of the “gold-standard” experiment while enhancing statistical efficiency.
A flexible approach for integrating randomized and observational data involves generating a biased estimate from the pooled data, learning a bias function between the data sources, and adjusting the biased estimate to obtain an unbiased causal effect estimate (Kallus et al., 2018; Wu and Yang, 2022; Shyr et al., 2023; van der Laan et al., 2024d). However, without model assumptions on the data-generating distribution, nonparametric efficient estimators like those constructed via debiased machine learning asymptotically gain no efficiency from including biased observational data (Dang et al., 2022). The efficiency gain from a larger sample size is completely offset by increased variance from learning the confounding bias function. When restricted to regular estimators, efficiency gains require imposing parametric or semiparametric restrictions on the bias function, which may induce estimation and confounding bias if these assumptions are violated. To address this limitation in cross-sectional studies, van der Laan et al. (2024d) proposed the ADML framework, which uses the highly adaptive lasso to learn model assumptions data-adaptively, yielding nonparametric superefficient estimators.
E.2 ADML Methodology
@@
As outlined in Section B, an ADML estimator can be constructed by learning a model $H_{n}$ and performing inference for the data-adaptive parameter $\Psi_{H_{n}}$. According to the theory in Section B, the ADML estimator is, under certain conditions, asymptotically linear and efficient for the oracle parameter $\Psi_{H_{0}}$ corresponding to a limiting oracle model $H_{0}$ to which $H_{n}$ converges. For example, $H_{n}$ can be learned using fitted Q-iteration for $q_{0}(a,1,s)$ with the Highly Adaptive Lasso (Benkeser and Van Der Laan, 2016), as in van der Laan et al. (2024d), where a preliminary estimate of the observational $Q$-function $q_{0}(a,0,s)$ is used as an offset, allowing the bias function to be directly modeled. Consequently, the ADML estimator is asymptotically linear and superefficient for $\Psi(P_{0})$ under the nonparametric model, maintaining robustness to model misspecification while gaining efficiency by pooling the two studies when there is learnable structure in the bias function. This approach directly generalizes the approach of van der Laan et al. (2024d) for short-term causal effects in cross-sectional studies, which corresponds to the case where $\gamma=0$.
We now propose a specific ADML estimator that leverages isotonic Bellman calibration, as described in Section 5, to construct superefficient estimators of $\Psi(P_{0})$ while avoiding the computational challenges and instability of estimating the Riesz representer via min-max optimization. Suppose the observational data is far larger than the experimental study, as often occurs in industrial applications where historical data vastly outnumbers randomized data from A/B tests. In this setting, the observational $Q$-function $q_{0}(a,0,s)$ can be estimated very accurately and is effectively known. Assuming $q_{0}(a,0,s)$ is known, we define the oracle model:
@@
To construct an ADML estimator, we use a modified version of calibrated FQI that incorporates $q_{0}$ as an offset. Define the modified outcome $\widetilde{Y}_{0}=Y_{0}+\gamma V_{q_{0}}^{\pi}(A_{1},0,S_{1})-q_{0}(A_{0},0,S_{0})$. By the Bellman equation for $q_{0}$, we have $E_{0}[\widetilde{Y}_{0}\mid A_{0},Z,\widetilde{S}_{0},Z=0]=0$ and $E_{0}[\widetilde{Y}_{0}\mid A_{0},S_{0},Z=1]=E_{0}[b^{\pi}_{0}(A_{0},Z,\widetilde{S}_{0})-\gamma V_{b^{\pi}_{0}}^{\pi}(1,S_{1})\mid A_{0},S_{0},Z=1]$, so $b^{\pi}_{0}$ satisfies the Bellman equation for $\widetilde{Y}_{0}$ given $Z=1$. In our modification of Algorithm 2, $q_{n}$ is replaced by $b_{n}^{\pi}$, each $Y_{0,i}$ is replaced by $\widetilde{Y}_{0,i}:=Y_{0,i}+\gamma V_{q_{0}}^{\pi}(0,S_{1,i})-q_{0}(A_{0,i},0,S_{0,i})$, and calibration is applied only using observations with $Z_{i}=1$. This corresponds to calibrated fitted Q-iteration where calibration uses the class $\{(a,s)\mapsto q_{0}(a,0,s)+zf(b_{n}(a,s)):f\text{ is an isotonic function}\}$. Applying this procedure, we obtain an isotonic-calibrated bias function $b_{n}^{*}$ that satisfies the empirical orthogonality condition for each transformation $f:\mathbb{R}\rightarrow\mathbb{R}$:
@@
A debiased plug-in estimator is then given by $\frac{1}{\sum_{i=1}^{n}Z_{0,i}}\sum_{i=1}^{n}Z_{0,i}m(A_{0,i},S_{0,i},q^{*}_{n,0})$, where $q^{*}_{n,0}(a,s)=q_{0}(a,0,s)+zb_{n}^{*}(a,s)$ is the calibrated estimator of $q_{0}^{(\pi)}$. An application of the results from Section 5 with the offset outcome $\widetilde{Y}_{0}:=Y_{0}+\gamma V_{q_{0}}^{\pi}(0,S_{1})-q_{0}(A_{0},0,S_{1,i})$ conditional on $Z_{0}=1$ establishes the asymptotic linearity and superefficiency of this estimator under the stated conditions.
Appendix F Proofs for Section 3
F.1 Proof of Theorem 8 on uniqueness of Bellman projection

### v4

@@ D.1 Simulation design
[…]
@@ D.2 Additional experimental results
Appendix E Code
The Python implementation of fitted $Q$-calibration is provided below.
Our implementation provides a scalable and efficient version of fitted $Q$-calibration using regression tree estimation via xgboost. We exploit the fact that isotonic regression is equivalent to fitting a monotone-constrained regression tree of unconstrained depth. The max_depth and min_child_weight parameters in isotonic regression can be specified or tuned without affecting calibration. We recommend setting max_depth to a large value, as the monotonicity constraint automatically limits effective depth. Setting min_child_weight to a small nonzero value (e.g., 30) helps mitigate the poor boundary behavior of standard isotonic regression (van2024stabilized).
import numpy as np
import xgboost
def isoreg_with_xgboost(x, y, max_depth=15, min_child_weight=20, weights=None):
    Fits isotonic regression using XGBoost with monotonic constraints to ensure
    non-decreasing predictions as the predictor variable increases.
    Args:
        x (np.array): A vector or matrix of predictor variables.
        y (np.array): A vector of response variables.
        max_depth (int, optional): Maximum depth of the trees in XGBoost.
                                   Default is 15.
        min_child_weight (float, optional): Minimum sum of instance weights
                                            needed in a child node. Default is 20.
        weights (np.array, optional): A vector of weights for each instance.
                                      If None, all instances are equally weighted.
    Returns:
        function: A prediction function that takes a new predictor variable x
                  and returns the model’s predicted values.
    # Create an XGBoost DMatrix object from the data with optional weights
    data = xgb.DMatrix(data=np.asarray(x).reshape(len(y), -1),
    label=np.asarray(y), weight=weights)
    # Set parameters for the monotonic XGBoost model
    params = {
        ’max_depth’: max_depth,
        ’min_child_weight’: min_child_weight,
        ’monotone_constraints’: "(1)",  # Enforce monotonic increase
        ’eta’: 1,
        ’gamma’: 0,
        ’lambda’: 0
    # Train the model with one boosting round
    iso_fit = xgb.train(params=params, dtrain=data, num_boost_round=1)
    # Prediction function for new data
    def predict_fn(x):
        Predicts output for new input data using the trained isotonic regression model.
        Args:
            x (np.array): New predictor variables as a vector or matrix.
        Returns:
            np.array: Predicted values.
        x = np.atleast_2d(x).T if x.ndim == 1 else x
        data_pred = xgb.DMatrix(data=x)
        pred = iso_fit.predict(data_pred)
        return pred
    return predict_fn
def fitted_q_calibration_iso(q_n, Y0, A0, S0, S1, gamma, V_pi, K=5, sample_weight=None):
    Fitted Q-Calibration with isotonic regression and early stopping.
    Parameters
    q_n : callable
        Base Q-estimator. Must support vectorized evaluation:
        ‘q_n(A0, S0)‘ -> array of shape (n,), and ‘q_n(a, s)‘ for new inputs.
    Y0 : array-like of shape (n,)
@@ Appendix E Code
    A0 : sequence length n
        Actions at time 0 used to compute base scores x_i = q_n(A0, S0).
    S0 : sequence length n
        States at time 0 corresponding to A0.
    S1 : sequence length n
@@
    gamma : float in [0, 1)
        Discount factor.
    V_pi : callable
        Value operator. Called as ‘V_pi(q_fn, S1)‘ and must return the vector
@@
        Here ‘q_fn(a, s)‘ is a callable Q-function supplied by this routine.
    K : int, default=5
        Maximum number of calibration iterations.
    sample_weight : array-like of shape (n,), optional
        Nonnegative weights for the isotonic fit. If None, uses uniform weights.
    Returns
    q_star : callable
[…]
        where K’ <= K due to early stopping.
    Y0 = np.asarray(Y0, float)
    n = len(Y0)
    eps = 1.0 / n
    sample_weight = np.ones(n) if sample_weight is None
    else np.asarray(sample_weight, float)
    # Fixed base scores x_i = q_n(A0, S0)
    x = q_n(A0, S0)
[…]
    q_star = q_n
    q_prev_vals = q_star(A0, S0)
    for _ in range(K):
        # Bellman targets with current q_star
        y = Y0 + gamma * V_pi(q_star, S1)
        # One-step isotonic fit via your monotone XGBoost helper
        iso = isoreg_with_xgboost(x, y,
        max_depth=15, min_child_weight=20, weights=sample_weight)
        # Compose: q_star(a, s) = iso( q_n(a, s) )
        q_star = (lambda iso=iso, q_n=q_n: (lambda a, s: iso(q_n(a, s))))()
        # Early stopping using empirical L2(P_n)
        q_new_vals = q_star(A0, S0)
        diff = np.sqrt(np.mean((q_new_vals - q_prev_vals) ** 2))
        if diff < eps:
            break
        q_prev_vals = q_new_vals
    return q_star
Appendix F DRL with sieve estimation and model selection
F.1 General Approach
Selecting an appropriate working model is challenging and can compromise inference due to model misspecification bias. Adaptive debiased machine learning (ADML) (van2023adaptive) provides a unified framework that combines debiased estimation with data-driven model selection to construct superefficient estimators of smooth functionals, adapting to the structure of the nuisance components. By learning model assumptions or feature representations directly from data, ADML facilitates valid inference while mitigating misspecification bias. The calibrated plug-in estimator introduced in the previous section is a special case of ADML. In this section, we extend the ADML framework to the MDP setting, showing how semiparametric DRL can be combined with model selection to construct estimators that adapt to the functional form of the $Q$-function $q_{0}$, going beyond calibration alone. We refer to this extension as Adaptive DRL (ADRL).
For example, the working model $H_{n}$ could be selected via cross-validated FQI over a sieve of models, i.e., a sequence of increasingly complex classes $H_{1}\subset H_{2}\subset H_{3}\subset\dots\subset H_{\infty}:=H$, where $H$ is a correctly specified model containing $q_{0}$. A plausible oracle submodel $H_{0}$ is the smallest correctly specified class in the sieve that contains $q_{0}$, which can feasibly be approximated via cross-validation. Alternatively, $H_{n}$ could result from a variable selection procedure or a learned feature transformation, with $H_{0}$ corresponding to the asymptotically selected variables or limiting transformation. Data-adaptive methods for learning state-action feature representations have been proposed in pritz2021jointly and pavse2024state. Such transformations can also be derived directly from the fitted FQI model $q_{n,H_{n}}^{\pi}$, for example, by one-hot encoding the leaf nodes of trees in a random forest or gradient-boosted tree model, as in Section 3.1 of he2014practical.
F.2 Asymptotic theory
Theorem 9 (Second-order model approximation bias).
where $H_{n,0}:=H_{n}\oplus H_{0}$ is the direct sum linear model.
To further clarify these conditions, suppose the working model $H_{n}:=H_{\phi_{n}}$ and the oracle model $H_{0}:=H_{\phi_{0}}$ are induced by feature transformations. For a transformation $\phi:\mathcal{A}\times\mathcal{Z}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define $H_{\phi}:=\{f\circ\phi:f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}$. The combined model $H_{n,0}$ is given by $H_{(\phi_{n},\phi_{0})}$, where $(\phi_{n},\phi_{0})$ denotes the feature map formed by stacking $\phi_{n}$ and $\phi_{0}$. Theorem 4 implies that the approximation bias vanishes if the nuisance functions derived from $\phi_{n}$ and $(\phi_{n},\phi_{0})$ converge to those derived from $\phi_{0}$. The special case $\phi_{n}=q_{n}$ and $\phi_{0}=q_{0}$ recovers Theorem 5 as a corollary. In Lemma 14 (Appendix J.1), we show that with features $X$ and outcome $Y$, the $L^{2}(P_{0})$ error of estimating $E_{0}[Y\mid\phi_{0}(X)]$ using either $E_{0}[Y\mid\phi_{n}(X),\mathcal{D}_{n}]$ or $E_{0}[Y\mid\phi_{n}(X),\phi_{0}(X),\mathcal{D}_{n}]$ is bounded by the feature approximation error $\sqrt{\int\|\phi_{n}(x)-\phi_{0}(x)\|_{\mathbb{R}^{m}}^{2}P_{0,X}(dx)}$. A sufficient condition for this bound is that the map $(t_{1},t_{2})\mapsto E_{0}[Y\mid\phi_{n}(X)=t_{1},\phi_{0}(X)=t_{2},\mathcal{D}_{n}]$ is almost surely Lipschitz continuous, generalizing Condition D2.
$\varphi_{P,H}(s,a,y,s^{\prime}):=\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{y+\gamma V^{\pi}(q_{P,H}^{\pi})(s^{\prime})-q_{P,H}^{\pi}(a,s)\right\}+m(s,a,q_{P,H})-\Psi_{H}(P).$
Let $\varphi_{n,H_{n}}$ denote the estimator of the influence function $\varphi_{0,H_{n}}$, obtained by plugging in our nuisance estimators.
  1. (C4)
  2. (C5)
    Nuisance estimation rate: $\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\|_{P_{0}}\|\mathcal{T}_{0}(q_{n,H_{n}})-\mathcal{T}_{0}(q_{0,H_{n}})\|_{P_{0}}=o_{p}(n^{-1/2})$.
  3. (C6)
  4. (C7)
Theorem 10.
Together, C4 and C5 imply that $\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})=(P_{n}-P_{0})\varphi_{0,H_{n}}+o_{p}(n^{-\frac{1}{2}})$, so that $\psi_{n,H_{n}}$ is debiased for the working parameter $\Psi_{H_{n}}(P_{0})$. Conditions C6 and C7, which ensures that data-driven model selection preserves the validity of the debiased machine learning estimator, appear in prior works on ADML (van2023adaptive; van2024adaptive). Condition C6 is an asymptotic stability condition requiring the EIF for the learned model $H_{n}$ to converge to the EIF for the oracle submodel $H_{0}$, which necessitates that $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ and $q_{0,H_{n}}$ are asymptotically consistent with their oracle counterparts. Condition C7 ensures the parameter approximation bias satisfies $\Psi_{n}(P_{0})-\Psi(P_{0})=o_{p}(n^{-\frac{1}{2}})$ in view of Theorem 4.
Appendix G Data-fusion application of ADML
G.1 Background on confounding-robust data-fusion
We consider a data fusion setting in which experimental data ($Z=1$) is augmented with historical control data ($Z=0$), adapting the frameworks of kallus2018removing and van2024adaptive to Markov decision processes. Define the state as $S_{t}=(Z,\widetilde{S}_{t})$, as in Example 2. Suppose the study indicator $Z$ denotes whether a study unit belongs to an observational study $(Z=0)$, such as historical data, or a randomized experiment $(Z=1)$, such as a randomized control trial. We address the data-fusion problem of augmenting experimental data with potentially biased observational data to increase power by effectively enlarging the sample size. While randomization ensures unbiasedness in the experimental study, incorporating observational data can introduce bias from unmeasured confounding unless strong, untestable assumptions are made. Our goal is to combine these data sources in a confounding-robust manner that retains the unbiasedness of the “gold-standard” experiment while enhancing statistical efficiency.
A flexible approach for integrating randomized and observational data involves generating a biased estimate from the pooled data, learning a bias function between the data sources, and adjusting the biased estimate to obtain an unbiased causal effect estimate (kallus2018removing; wu2022integrative; shyr2023multi; van2024adaptive). However, without model assumptions on the data-generating distribution, nonparametric efficient estimators like those constructed via debiased machine learning asymptotically gain no efficiency from including biased observational data (dang2022cross). The efficiency gain from a larger sample size is completely offset by increased variance from learning the confounding bias function. When restricted to regular estimators, efficiency gains require imposing parametric or semiparametric restrictions on the bias function, which may induce estimation and confounding bias if these assumptions are violated. To address this limitation in cross-sectional studies, van2024adaptive proposed the ADML framework, which uses the highly adaptive lasso to learn model assumptions data-adaptively, yielding nonparametric superefficient estimators.
G.2 ADML Methodology
In this section, we extend the ADML framework of van2024adaptive to nonparametric data fusion in MDPs. Following the experimental grounding approach of kallus2018removing, we define the confounding bias function as the difference in $Q$-functions, $b_{0}^{\pi}(a,s)=q_{0}(a,1,s)-q_{0}(a,0,s)$, between the experimental and observational studies. Our parameter of interest is defined as $\Psi(P)=E_{P}[m(A_{0},Z,\widetilde{S}_{0},q_{P})]$, where the mapping $m(s,a,q)\mapsto\int q_{P}(a^{\prime},s)\pi(a^{\prime}\mid s)dz$ represents the long-term effect of policy $\pi$. This effect is averaged over the study-pooled covariate distribution; alternatively, the covariate distribution could be defined over the experimental study—see van2024adaptive for details. We approximate the parameter $\Psi$ by $\Psi_{H}=E_{P}[m(A_{0},Z,\widetilde{S}_{0},q_{P,H})]$, where $H$ is the partially linear working model that imposes the semiparametric restriftion that the bias function $b_{0}^{\pi}$ lies in a Hilbert space $\mathcal{B}\subseteq L^{2}(P_{A_{0},Z,\widetilde{S}_{0}})$. For example, one may posit that $b_{0}^{\pi}$ is well approximated by linear combinations of some finite set of features derived from $(A_{0},Z,\widetilde{S}_{0})$.
As outlined in Section F, an ADML estimator can be constructed by learning a model $H_{n}$ and performing inference for the data-adaptive parameter $\Psi_{H_{n}}$. According to the theory in Section F, the ADML estimator is, under certain conditions, asymptotically linear and efficient for the oracle parameter $\Psi_{H_{0}}$ corresponding to a limiting oracle model $H_{0}$ to which $H_{n}$ converges. For example, $H_{n}$ can be learned using fitted Q-iteration for $q_{0}(a,1,s)$ with the Highly Adaptive Lasso (benkeser2016highly), as in van2024adaptive, where a preliminary estimate of the observational $Q$-function $q_{0}(a,0,s)$ is used as an offset, allowing the bias function to be directly modeled. Consequently, the ADML estimator is asymptotically linear and superefficient for $\Psi(P_{0})$ under the nonparametric model, maintaining robustness to model misspecification while gaining efficiency by pooling the two studies when there is learnable structure in the bias function. This approach directly generalizes the approach of van2024adaptive for short-term causal effects in cross-sectional studies, which corresponds to the case where $\gamma=0$.
We now propose a specific ADML estimator that leverages fitted $Q$-calibration, as described in Section 4, to construct superefficient estimators of $\Psi(P_{0})$ while avoiding the computational challenges and instability of estimating the Riesz representer via min-max optimization. Suppose the observational data is far larger than the experimental study, as often occurs in industrial applications where historical data vastly outnumbers randomized data from A/B tests. In this setting, the observational $Q$-function $q_{0}(a,0,s)$ can be estimated very accurately and is effectively known. Assuming $q_{0}(a,0,s)$ is known, we define the oracle model:
$H_{0}:=\left\{q_{0}(a,0,s)+f(b_{0}^{\pi}(a,s)):f\text{ is a real-valued transformation}\right\}.$
This model consists of all $Q$-functions that agree with $q_{0}$ on the observational study and have a bias function differing from $b_{0}^{\pi}=q_{0}(a,1,s)-q_{0}(a,0,s)$ by a transformation of an arbitrary one-dimensional function. Notably, this model is necessarily correctly specified, as taking $f$ to be the identity function recovers $q_{0}$. Given an initial estimator $b_{n}^{(\pi)}$ of $b_{0}^{(\pi)}$, we approximate the oracle model by the working model: $H_{n}:=\left\{q_{0}(a,0,s)+f(b_{n}^{\pi}(a,s)):f\text{ is a real-valued transformation}\right\},$ and propose to obtain superefficient inference for $\Psi(P_{0})$ by constructing ADML estimators based on $\Psi_{H_{n}}$.
@@ G.2 ADML Methodology
To construct an ADML estimator, we use a modified version of calibrated FQI that incorporates $q_{0}$ as an offset. Define the modified outcome $\widetilde{Y}_{0}=Y_{0}+\gamma V^{\pi}(q_{0})(A_{1},0,S_{1})-q_{0}(A_{0},0,S_{0})$. By the Bellman equation for $q_{0}$, we have $E_{0}[\widetilde{Y}_{0}\mid A_{0},Z,\widetilde{S}_{0},Z=0]=0$ and $E_{0}[\widetilde{Y}_{0}\mid A_{0},S_{0},Z=1]=E_{0}[b^{\pi}_{0}(A_{0},Z,\widetilde{S}_{0})-\gamma V_{b^{\pi}_{0}}^{\pi}(1,S_{1})\mid A_{0},S_{0},Z=1]$, so $b^{\pi}_{0}$ satisfies the Bellman equation for $\widetilde{Y}_{0}$ given $Z=1$. In our modification of Algorithm 1, $q_{n}$ is replaced by $b_{n}^{\pi}$, each $Y_{0,i}$ is replaced by $\widetilde{Y}_{0,i}:=Y_{0,i}+\gamma V^{\pi}(q_{0})(0,S_{1,i})-q_{0}(A_{0,i},0,S_{0,i})$, and calibration is applied only using observations with $Z_{i}=1$. This corresponds to calibrated fitted Q-iteration where calibration uses the class $\{(a,s)\mapsto q_{0}(a,0,s)+zf(b_{n}(a,s)):f\text{ is an isotonic function}\}$. Applying this procedure, we obtain an isotonic-calibrated bias function $b_{n}^{*}$ that satisfies the empirical orthogonality condition for each transformation $f:\mathbb{R}\rightarrow\mathbb{R}$:
@@
A debiased plug-in estimator is then given by $\frac{1}{\sum_{i=1}^{n}Z_{0,i}}\sum_{i=1}^{n}Z_{0,i}m(A_{0,i},S_{0,i},q^{*}_{n,0})$, where $q^{*}_{n,0}(a,s)=q_{0}(a,0,s)+zb_{n}^{*}(a,s)$ is the calibrated estimator of $q_{0}^{(\pi)}$. An application of the results from Section 4 with the offset outcome $\widetilde{Y}_{0}:=Y_{0}+\gamma V^{\pi}(q_{0})(0,S_{1})-q_{0}(A_{0},0,S_{1,i})$ conditional on $Z_{0}=1$ establishes the asymptotic linearity and superefficiency of this estimator under the stated conditions.
Appendix H Proofs for Section 3
H.1 Proof of Theorem 7 on uniqueness of Bellman projection